7:30: power up of DX and SX systems. RR cubes are mounted back on both sides.

Side PCT
SX 20160310_151257_pcTransmission.jpg
DX 20160310_155307_pcTransmission.jpg

TipTiltSystemSx calibrations

We setup the SX AGW to be used with the internal source, FW1 has 90-10, FW2 QC mirror.

Board -73.9 -50.0 0.0
Then we center the QC putting the Z stage at 25mm, so 1" spot:

Board -72.3 -49.8 25.0
We change FW1 to dichroic 600-1000nm:

Board -73.05 -50.1 25.0
We find that a better position for the cube rotator is 91.45deg instead of the 92.2deg of the board setup.

IM REC
20160310_180407 20160310_181500
20160310_182100  
These IMs are very different from the one recorded in December, we have almost a factor 100 stronger signals... Could be due to the use of different pyramid modulations, since the QC signals are re-normalized on Truth sensing modes to get the proper orientation wrt the ASM tip and tilt modes. So we change the SW, now thruth modes are used just to identify the direction of the modes, not their amplitude.

New board position for Z=35mm, in focus on IS:

Board -72.6 -50.3 35.0

IM REC
20160310_191345 20160310_191400
20160310_195553 wrong IM
20160310_201326  
20160310_202717 20160310_203600
We build a combined rec from the last QC rec:

  • COMBO REC (TCoG48 + QC): 20160310_204000

That REC was acquired with LUCI rotator at 92deg.

We test in close loop using the ARGOS CalUnit. We found a problem of gain: the REC seems a factor 100 too big! We build a new combo REC dividing the 20160310_223600 by a factor 100:

IM REC
20160310_202717 20160310_213500

  • COMBO REC (TCoG48 + QC): 20160310_213600

We test it in closed loop and it seems to work...

TipTiltSystemSx co-pointing with Pyramid WFS

We start measuring the hot spot on CCD47 for QC:

Bin FW1 Hot pixel
4 90-10 [29.4; 40.4]
White light source is set to 65 steps. This is equivalent to a mag 13.5 on PYR.

We spin the LUCI rotator while we keep closed the TT loop on QC. We track with the pupil derotator and we find the LUCI rotator angle that minimizes the pointing difference between the 2 sensors. This value is 65deg.

We measure the hot spot on CCD47 when the spot is centred on PYR: [32.6; 40.4].

A possible procedure to co-point the 2 sensors at a given rotator angle is:

  • close the TT loop on PYR
  • spin FW2 to minimize the QC signals (= difference in 4 channel fluxes)
  • offset the pyramid modulator to null the QC signals
  • close TT loop on QC and measure the residual TT from truth modes

We keep the TT loop closed on QC and we measure the piezo offset needed to null the pointing difference:

Rot angle X offset Y offset
65 2.2 1.3
55 2.2 1.4
0 3.3 2.7
-35 4.5 2.1
115 2.4 -1.2
155 3.3 -2.5
180 4.2 -2.8
210 5.0 -2.0
We then try to move FW2 and to minimize the pointing offset at 245deg.

FW2 X offset Y offset
-0.003 >5.0 ...
0.003 4.2 -1.3
0.005 3.4 -0.2

LgswUnitDx to CalUnit alignment

In parallel to the SX QC calibrations we try to align the CalUnit to LGSW on DX side. First alignment was done yesterday when Gustavo was looking at the Entrance windows flange of the sensor. Today we refined this position trying to reduce the overall coma. We ended up to a limit position were the decenter Y of the hexapod was hitting the limit:

DX Hexa 4.80 -8.45 -2.00 50 -21
At the end of daytime WG moved the upper arm of the CalUnit gusset by ~1mm with the aim to reduce the hexapod decenter Y.

On top of the alignment work we tryed also to acquire a 10 modes LGSW IM, but just one LGS at time could be feeded inside the WFS. We build a just for the Blue pupil using the old KL_v16 as modal basis:

IM REC COMBO
20160310_193215 20160310_193700 20160310_193900
We tryed also to close the loop using this REC but the tilt was diverging...

-- MarcoBonaglia - 11 Mar 2016
Topic revision: r1 - 11 Mar 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback